OpenStudy (anonymous):

How do I convert 41in^2 to cm^2?

OpenStudy (mathmale):

What is the relevant conversion factor? I have this impressed in memory: 1 inch = 2.54 cm. You could start with 41 in^2, Enclose that "in" inside partentheses: 41 (in)^2, and then replace "in" with "2.54 cm". Willing to try that?

OpenStudy (anonymous):

\[41inch ^{2}=41*2.54*2.54 cm ^{2}=?\]

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} in&cm\\ \hline\\ 1&2.54\\ 1in^2& 2.54^2\\ \hline\\ 41in^2&41( 2.54)^2\implies 41\cdot 2.54^2 \end{array}\)

OpenStudy (mathmale):

JD: Apprreciate your acting on my suggestion. My only additional suggestion is that your final result should have units of measurement as an integral part.

OpenStudy (jdoe0001):

ohh yeah, ahemm =)

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} 1in&2.54cm\\ 1in^2& (2.54cm)^2\implies 2.54^2cm^2\\ \hline\\ 41in^2&41( 2.54^2)cm^2 \end{array}\)